Hello,
Could you give me the surface area of a cylinder when the radius is 6cm and the Length is 21cm.
Could you show me the full step by step guide to work this out.
Many thanks

Found 2 solutions by nycsub_teacher, ankor@dixie-net.com:
Answer by nycsub_teacher(90)

Could you give me the surface area of a cylinder when the radius is 6cm and the Length is 21cm.
You need a formula and it is
surface area = 2(πr^2) + 2(πr)(h), where r = radius, h = height
and π = pi = 3.14
I guess by length you meant to type height.
Let SA = surface area of cynlinder for short.
SA = 2(3.14)(6)^2 + 2(3.14)(21)
SA = 226.08 + 131.88
SA = 357.96cm^2

Answer by ankor@dixie-net.com(22740)

Could you give me the surface area of a cylinder when the radius is 6cm and the Length is 21cm.
:
The surface area of cylinder consists of the parts
Part 1: the height times the circumference (in this case it's the length)
Part 2: the area of the two ends
;
SA = (2*pi*r*h) + (2*pi*r^2)
:
In your problem; r=6, h=21
SA = (2*pi*6*21) + (2*pi*6^2)
;
SA = 791.68 + 266.19
:
SA = 1017.87 sq/cm